OpenStudy (anonymous):

Which mathematical sentence represents the solution for a in the equation 5b = bc + 4a?

OpenStudy (anonymous):

\[a = \frac{5b - bc}{4} \implies a = \frac{b(5-c)}{4}\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

\[\huge \color{green}{\textbf{Welcome To Openstudy as Well..}}\]

OpenStudy (anonymous):

Thank you very much